OpenStudy (anonymous):

Nathan wrote the following steps to prove the equation: Given: 3(x + 2) = 6x + 3 Prove: x = 1 Step Mathematical Statement Justification 0 3(x + 2) = 6x + 3 Given 1 3x + 6 = 6x + 3 Distributive Property 2 6 = 3x + 3 Commutative Property of Addition 3 3x + 3 = 6 Symmetric Property of Equality 4 3x = 3 Subtraction Property of Equality 5 x =1 Division Property of Equality Which step of justification is incorrect, and what should the justification for that step be to solve the equation?

OpenStudy (anonymous):

The answers: Step 4; Commutative Property of Multiplication Step 1; Associative Property of Addition Step 2; Subtraction Property of Equality Step 3; Reflexive Property of Equality

OpenStudy (anonymous):

Could you help me?

OpenStudy (anonymous):

the key is knowing all of those properties... How is step 2 obtained from step 1? Definitely not commutative propert....

OpenStudy (anonymous):

you have to do addition before subtraction right?

OpenStudy (anonymous):

Leaving doesn't help me learn, can someone else please help me?

OpenStudy (anonymous):

What else dpalnc, was I right about that?

OpenStudy (anonymous):

about leaving?

OpenStudy (anonymous):

I was wrong, but it had said you left, I apologize for my false assumptions.

OpenStudy (anonymous):

Can you still help me?

OpenStudy (anonymous):

step1: 3x + 6 = 6x + 3 step2: 6 = 3x + 3 notice to get from step 1 to step 2 you'd have to subtract 3x from both sides of the equal sign... so 3x + 6 -3x = 6x + 3 -3x 6 = 3x + 3 this is what step 2 is.... so the property used is subtraction property of equality... just a fancy way of saying "subtract the same amount on both sides of the equal sign"

OpenStudy (anonymous):

So the answer is 3 then?

OpenStudy (anonymous):

the third choice on your list...

OpenStudy (anonymous):

I mean the third answer of the four given.

OpenStudy (anonymous):

yes, would you like to help me with one more now that I understand that one?

OpenStudy (anonymous):

sure... just post it up and close this one... :)

OpenStudy (anonymous):

alirght